f(x+h)-f(x) Find if f(x)=3x2 + 5x-18 h O A. 6x+5-3h 0 B. 6x +5+h 0 C. 6x+5 O D. 6x + 5+ 3h
Answer #1

Given fini = 3x² +51-18 finthi = 3(xth)? +5(a+h) – 18 = 31x²th²+zahl +50+5h-18 = 3x² + 3h² +6x2 + 5x+5h-18 futhi feal 17 n 3n